suppose that the width of a rectangle is 5 inches shorter than the length and that the perimeter of the rectangle is 50. set up an equation for the perimeter involving L, the length of the rectangle. then solve algebraically.
Answer by checkley75(3666)

W=L-5
2(L-5)+2L=50
2L-10+2L=50
4L=50+10
4L=60
L=60/4
L=15 ANSWER FOR THE LENGTH.
W=15-5
W-10 ANSWER FOR THE WIDTH.
PROOF
2*15+2*10=50
30+20=50
50=50